OpenStudy (anonymous):

A package of pencils cost $6.25 per set at an office supply store. Which equation can be used to determine the price, P, of n sets of these pencils? n = 6.25P P = 6.25n n = 6.25/ P P = 6.25/ n

Mehek (mehek14):

which one do you think?

OpenStudy (anonymous):

a

Mehek (mehek14):

correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

one more?

OpenStudy (anonymous):

I can help with your next question

OpenStudy (anonymous):

Marneshia walked three over eight of a mile in three over five of an hour. What equation can be used to calculate her unit rate in miles per hour? 3/8 over 3/5 = 5/8 mile per hour 3/8 over 3/5 = 1 and 3/5 miles per hour 3/5 over 3/8 = 1 and 3/5 miles per hour 3/5 over 3/8= 5/8 mile per hour

Mehek (mehek14):

A

OpenStudy (anonymous):

how

OpenStudy (anonymous):

yeah A

OpenStudy (anonymous):

you do 3/8*5/3

Mehek (mehek14):

because 3/8 / 3/5 = 3/8= 5/3 = 5/8

OpenStudy (anonymous):

you get 15/24 then you must reduce

OpenStudy (anonymous):

to get A

Mehek (mehek14):

or just cancel out the 3's and you ave 5/8

OpenStudy (anonymous):

i dont know who to give a medal to now :/

OpenStudy (anonymous):

i dont know who to give a medal to now :/